JA buffer is prepared by dissolving 3.58 g of for c acid (????) and 5.95 g of potassium) formate (KCOH) in enough water to make 250 mL. of solution. What is the pH? Formic acid has a pK, of 3.80.
Answer #1

First we must calculate the moles of salt and acid present, and then its concentration in the final solution:

nHCO2H = 3.58 g * (1 mol / 46.03 g) = 0.0778 mol

nKCO2H = 5.95 g * (1mol / 84.11 g) = 0.0707 mol

[HCO2H] = 0.0778 mol / 0.25 l = 0.3112 M

[KCO2H] = 0.0707 mol / 0.25 l = 0.2828 M

Finally we calculate the pH with the help of the Henderson-Hasselbach equation:

pH = pKa + Log ([KCO2H] / [HCO2H]) = 3.8 + Log (0.2828 / 0.3112) = 3.76.

pH = 3.76.